Outline Getting Started

Cluedump – L^{A}TEX

Jason Gross — jgross@mit.edu

November 2, 2010

Outline

Getting Started Installing L^{A}TEX What is L^{A}TEX?

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

L^{A}TEX on Linux

▶ Usually comes preinstalled ▶ TEXLive can be downloaded from

  http://www.tug.org/texlive/

▶ Use your favorite text editor (vim, emacs, etc.), OR ▶ Specialized editors for L^{A}TEX, e.g. kile

  (http://kile.sourceforge.net/)
    ▶ sudo apt-get install kile okular

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

L^{A}TEX on Windows

▶ proTEXt installs MiKTEX and TEXnicCenter

▶ Download from http://www.tug.org/protext/

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

L^{A}TEX on Mac

▶ MacTEX (http://www.tug.org/mactex/)
▶ TEXShop (http://pages.uoregon.edu/koch/texshop/)

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

What L^{A}TEX is

▶ A typesetting system ▶ Aimed at math and text ▶ Extensible ▶ A macro-based Turing complete programming language

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

What L^{A}TEX is not

▶ A WYSIWYG editor

▶ A programming language

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

What L^{A}TEX is not

▶ A WYSIWYG editor

▶ A programming language

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

What L^{A}TEX is not

▶ A WYSIWYG editor

▶ A programming language

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

What L^{A}TEX is not

▶ A WYSIWYG editor

▶ A programming language

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

Guiding Principles

▶ You’re not a professional typesetter! Don’t override L^{A}TEX’s default formatting unless you have a very good reason for doing so.

▶

Outline Getting Started Installing L^{A}TEX What is L^{A}TEX?

Guiding Principles

▶ You’re not a professional typesetter! Don’t override L^{A}TEX’s default formatting unless you have a very good reason for doing so.

▶